In a certain food stall, 278 out of 500 randomly selected consumers to indicate their preference for a new kind of food combination. Use a 99% confidence interval to estimate the true proportion ρ who like the new food combination.


1
Expert's answer
2022-05-05T16:48:09-0400

The sample proportion is computed as follows, based on the sample size "N = 500" and the number of favorable cases "X = 278:"


"\hat{p}=\dfrac{X}{N}=\dfrac{278}{500}=0.556"

The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}})"

"=(0.556-2.5758\sqrt{\dfrac{0.556(1-0.556)}{500}},"

"0.556+2.5758\sqrt{\dfrac{0.556{p}(1-0.556)}{500}})"

"=(0.499, 0.613)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.499<p<0.613," which indicates that we are 99% confident that the true population proportion "p" is contained by the interval "(0.499, 0.613)."